# The Conjugate Function Method for Surfaces with Elaborate Topological Types

### v1

@@ 5 Torii
Even though the conjugate function method is not directly applicable on surfaces of genus 1, if the parameter domain can be decomposed into admissible parts, then the conformal map can be constructed by considering each part independently.
@@ 5.1 Torus: Genus 1
For a regular torus the parameter domain is simply $0,2\pi]\times[0,2\pi]$. Due to symmetry, one admissible partition is to divide the parameter domain into four parts. The resulting map and the convergence graph are shown in Figure 4. The parameterisation of the torus, with inner radius $r=1/2$, and outer $R=3/2$ is
@@
Indeed, with a suitable reparameterisation of $v$ one can find an isothermal parameterisation for a torus.
@@ 5.2 Two examples: Genus 2 and Genus 3
and Euler’s relation $\chi=V-E+F$, one obtains
@@
so we need for a valid subdivision of the surface, for example, eight pentagons. By close inspection of the checkerboard image of the 2-torus, one can see that there is, indeed, one pentagon per every symmetric section, and hence, we conclude that the quasiconformal map is valid.
@@ 6.1 Case A: Nasser’s Challenge
The convergence graphs are shown in Figure 7c. Since the mesh (see Figure 7b) is generated on the original boundary discretisation, we cannot expect exponential convergence. Indeed, even though the reciprocal error is relatively small, the observed rate is only algebraic. The auxiliary subspace estimates for the primary and conjugate problems are optimistic. Notice that since the reciprocal error measures a product, it should stay above the auxiliary subspace estimates.
@@
Figure 7: Nasser’s Configuration. (a) Canonical domain. (b) Mesh used in computations. (c) Convergence graphs for error estimates. NN is the number of degrees of freedom. The data points represent the results on a fixed mesh with the constant polynomial order ranging from $p=2,\ldots,10$. Solid line: Reciprocal error; Dashed and dotted: Auxiliary subspace estimates for the primary and conjugate problems, respectively (loglog-plot).
@@ 6.2 Case B: Random Segments Inside a Rectangle
The resulting map is shown in Figure 8. The detail shown in Figure 8b shown how the slit perturbs the ”flow lines” locally. The slit canonical domain is illustrated in Figure 9a. Together with the map on the domain (see Figure 8a) one can trace the paths from edge to edge where the gridlines do or do not touch any of the slits.
@@
Figure 9: Random Segments. (a) Canonical slit domain. (b) Convergence graphs for error estimates. NN is the number of degrees of freedom. The data points represent the results on a fixed mesh with the constant polynomial order ranging from $p=2,\ldots,10$. Solid line: Reciprocal error; Dashed and dotted: Auxiliary subspace estimates for the primary and conjugate problems, respectively (loglog-plot).
@@ 8 Conclusions
In this work we have generalised and refined the conjugate function method to achieve the same level of accuracy on simply and multiply connected planar domains and Riemann surfaces. In contrast to the Koebe iteration, the construction presented here can be viewed as direct, with predictable a priori computational complexity. Our implementation relies on high-order finite element methods, both for obtaining high accuracy and the construction of the conjugate problem. This does not mean that the key observation in the minimisation process depends on one particular method for solving the underlying partial differential equations.
Extension to high-genus surfaces remains a future challenge. Automatic detection of charts for an atlas is a very difficult problem and we have not made any attempts to address this question. In cases where the symmetries suggest parameterisations, such as the 2-torus and 3-torus considered above, the conjugate function method can be used to compute highly accurate quasiconformal maps.
@@ Acknowledgments
A. Rasila and Y. Zheng were supported by NSF of Guangdong Province (No. 2024A1515010467) and Li Ka Shing Foundation GTIIT-STU Joint Research Grant (No. 2024LKSFG06)

### v2

@@ 5 Torii
The conjugate function method is not directly applicable on closed surfaces of genus 1 or higher. However, if the parameter domain can be decomposed into admissible parts, then the conformal map can be constructed by considering each part independently.
@@ 5.1 Torus: Genus 1
For a regular torus the parameter domain is simply $0,2\pi]\times[0,2\pi]$. Due to symmetry, one admissible partition is to divide the parameter domain into four parts. The resulting map and the convergence graph are shown in Figure 4. The parameterisation of the torus, with inner radius $r=1/2$, and outer radius $R=3/2$ is
@@
Indeed, with a suitable parameterization of $v$ one can find an isothermal parameterisation for a torus.
@@ 5.2 Two examples: Genus 2 and Genus 3
and Euler’s relation states that $\chi=V-E+F$, one obtains
@@
so we need for a valid subdivision of the surface, for example, eight pentagons. By close inspection of the checkerboard image of the 2-torus, one can see that there is, indeed, one pentagon per every symmetric section, and hence, we conclude that the construction is valid.
@@ 6.1 Case A: Nasser’s Challenge
The convergence graphs are shown in Figure 7c. Since the mesh (see Figure 7b) is generated on the original boundary discretisation, we cannot expect exponential convergence. Indeed, even though the reciprocal error is relatively small, the observed rate is only algebraic. The auxiliary subspace estimates for the primary and conjugate problems are optimistic. Notice that since the reciprocal error measures a product, it should stay above the auxiliary subspace estimates of both the primary and the conjugate problems.
@@
Figure 7: Nasser’s Configuration. (a) Canonical domain. (b) Mesh used in computations. (c) Convergence graphs for error estimates. NN is the number of degrees of freedom that depends on the mesh and the polynomial order. The data points represent the results on a fixed mesh with the constant polynomial order ranging from $p=2,\ldots,10$. Solid line: Reciprocal error; Dashed and dotted: Auxiliary subspace estimates for the primary and conjugate problems, respectively (loglog-plot).
@@ 6.2 Case B: Random Segments Inside a Rectangle
The resulting map is shown in Figure 8. The detail shown in Figure 8b shown how the slit perturbs the “flow lines” locally. The slit canonical domain is illustrated in Figure 9a. Together with the map on the domain (see Figure 8a) one can trace the paths from edge to edge where the gridlines do or do not touch any of the slits.
@@
Figure 9: Random Segments. (a) Canonical slit domain. (b) Convergence graphs for error estimates. NN is the number of degrees of freedom that depends on the mesh and the polynomial order. The data points represent the results on a fixed mesh with the constant polynomial order ranging from $p=2,\ldots,10$. Solid line: Reciprocal error; Dashed and dotted: Auxiliary subspace estimates for the primary and conjugate problems, respectively (loglog-plot).
@@ 7 Application: Texture Mapping
8 Polyhedral Surfaces
We conclude the set of examples with a polyhedral closed surface [28]. This is illustrated in Figure 15. The original geometric model is triangulated into 5026 3D triangles. The primary and conjugate problems are set simply by dividing the bottom rim into four sections. The image clearly shows the crowding effect, the sizes of the images of the cells of the uniform grid on the canonical are different on different parts of the domain. Since there is no global parameterization of the surface, in other words, no analytic chart is available, the variational formulation is adjusted on every surface triangle.
Figure 15: Polyhedral surface. The geometric model on the left is triangulated into 5026 3D triangles. Using the bottom of the statue as the boundary, a conformal map of the surface is computed.
@@ 9 Conclusions
In this work we have generalized and refined the conjugate function method to achieve the same level of accuracy on simply and multiply connected planar domains and Riemann surfaces. In contrast to the Koebe iteration, the construction presented here can be viewed as direct, with predictable a priori computational complexity. Our implementation relies on high-order finite element methods, both for obtaining high accuracy and the construction of the conjugate problem. This does not mean that the key observation in the minimisation process depends on one particular method for solving the underlying partial differential equations.
Extension to closed high-genus surfaces remains a future challenge. Automatic detection of charts for an atlas is a very difficult problem and we have not made any attempts to address this question. In cases where the symmetries suggest parameterisations, such as the 2-torus and 3-torus considered above, the conjugate function method can be used to compute highly accurate quasiconformal maps.
@@ Acknowledgments
A. Rasila and Y. Zheng were supported by NSF of Guangdong Province (No. 2024A1515010467), Li Ka Shing Foundation GTIIT-STU Joint Research Grant (No. 2024LKSFG06), and GTIIT Education Foundation.